# Is Sarcasm Detection A Step-by-Step Reasoning Process in Large Language Models?

### v1

@@ 3.3 Graph of Cues
The linear structure of CoC restricts it to a single path of reasoning. To fill this gap, we introduce graph of cues, a GoT-style paradigm that allows LLMs to flexibly choose and weigh multiple cues, unconstrained by the need for unique predecessor nodes (Fig. 2 (b)). GoC frames the problem of sarcasm detection as a search over a graph, and is formulated as a tuple $\left(\mathcal{M},\mathcal{G},\mathcal{E}\right)$, where $\mathcal{M}$ is the cue maker used to define what are the common cues, $\mathcal{G}$ is a graph of “sarcasm detection process”, $\mathcal{E}$ is cue evaluator used to determine which cues to keep selecting and in which order. Unlike ToT and GoT, GoC does not involve the modules of “thought generator” and “thought aggregation”.
@@
1. Cue maker. Human sarcasm judgment often relies on the combination and analysis of one or more cues to achieve an accurate understanding. Such cues can be broadly categorized into three types: linguistic cues, contextual cues and emotional cues. Linguistic cues refer to the linguistic features inherent in the text, including keywords, rhetorical devices, punctuation and language style. Contextual cues refer to the environment and background of the text, including topic, cultural background, common knowledge. Emotional cues denote the emotional stance conveyed by the text, including emotional words, special symbols (such as emojis) and emotional contrasts. A total number of 4+3+3=10 cues are adopted.
@@
2. Graph construction. In $\mathcal{G}=\left(V,E\right)$, the cues are regarded as vertices constituting the vertex set $V$, while the relations across cues form the edge set $E$. If there is an edge between cues $c_{k}$ and $c_{j}$, it is considered that $c_{k}$ and $c_{j}$ are closely related. Given the cue $c_{k}$, the cue evaluator $\mathcal{E}$ considers cue $c_{j}$ to provide the most complementary information to $c_{k}$, which would combine with $c_{k}$ to facilitate a deep understanding of sarcasm.
@@
3. Cue evaluator. We involve $\mathcal{G}$ in the LLM detecting sarcasm process. To advance this process, the cue evaluator $\mathcal{E}$ assesses the current progress towards judging sarcasm by means of determining whether the cumulative cues obtained so far are sufficient to yield an accurate judgment. If so, the search goes to an end. Otherwise, it serves as a heuristic for the search algorithm, determining which additional cues to select and in what order, to further the detection process. Similar to ToT, an LLM is used as the cue evaluator $\mathcal{E}$.
@@
We employ a voting strategy to determine the most valuable cue for selection, by explicitly comparing multiple potential cue candidates in a voting prompt, such as:
@@
Given an input text $\mathcal{X}$, the target is to accurately detect sarcasm. Now, we have collected the keyword information as the first step: {keywords}, judge if this provides over 95% confidence for accurate detection. If so, output the result. Otherwise, from the remaining cues {rhetorical devices, punctuation, …}, vote the most valuable one to improve accuracy and confidence for the next step. This step can be formulated as $\mathcal{E}\left(\mathcal{L}_{\theta}^{GoC},c_{j+1}\right)\sim Vote\left\{\mathcal{L}_{\theta}^{GoC}\left(c_{j+1}|\mathcal{X},c_{1,2,...,j}\right)\right\}_{c_{j+1}\in\{c_{j+1},...,c_{k}\}}$. In a nutshell, it greedily selects the most valuable cue until the final judgment is reached.
@@
Although the GoC enables the exploration of many possible paths across the cue graph, its nature remains grounded in a step-by-step reasoning paradigm (see Algorithm 2 in App. A).
@@ 3.4 Bagging of Cues
We further relax the assumption that the cues for sarcasm detection are inter-related. We introduce bagging of cues, an ensemble learning based paradigm that allows LLMs to independently consider varied combinations of cues without assuming a fixed order or dependency among them (Fig. 2 (c)).
BoC constructs a pool of the pre-defined $k=10$ cues $\mathcal{C}$. From this pool, $\mathcal{T}$ subsets are random sampled, each consisting of $q~{}\left(i.e.,1\leq q\leq k\right)$ cues. BoC thus leverages LLMs to generate $\mathcal{T}$ independent sarcasm predictions $\hat{y}_{t}$ based on the cues of each subset. Finally, such predictions are aggregated using a majority voting mechanism to produce the final sarcasm detection result. This approach embraces randomness in cue selection, enhancing the LLM’s ability to explore numerous potential paths, thus improving the robustness and accuracy of sarcasm detection. BoC consists of the following key steps:
Step 1. Cue subsets construction. A total of $\mathcal{T}$ cue subsets $\mathcal{S}_{t\in[1,2,...,\mathcal{T}]}=\left\{(c_{t_{1}},c_{t_{2}},...,c_{t_{q}}),t\in[1,2,...,\mathcal{T}]\right\}$ are created by randomly sampling without replacement from the complete pool of cues $\mathcal{C}$. Each sampling is independent.
Step 2. LLM prediction. For each subset $\mathcal{S}_{t}$, an LLM $\mathcal{L}_{\theta}^{BoC}$ is used to independently make sarcasm prediction through the comprehensive analysis of the cues in the subset and the input text. This can be conceptually encapsulated as $\hat{y}_{t}\sim\mathcal{L}_{\theta}^{BoC}\left(\mathcal{Y}|\mathcal{S}_{t},\mathcal{X}\right)$.
Step 3. Prediction aggregation. These individual predictions are then combined using an aggregation function, i.e., majority voting, to yield the final prediction: $Y\sim Vote\left(\left\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{\mathcal{T}}\right\}\right)$.
BoC treats all cues as independent and does not follow the step-by-step reasoning paradigm for sarcasm detection (see Algorithm 3 in App. A).
@@ 3.5 Tensor of Cues
CoC and GoC methods mainly handle low-order interactions between cues, while BoC assumes cues are independent. To capture high-order interactions among cues, we introduce tensor of cues, a novel paradigm that allows LLMs to amalgamate three types of cues ($viz.$ liguistic, contextual and emotional cues) into a high-dimensional representation (Fig. 2 (d)).
ToC treats each type of cues as an independent, orthogonal view for sarcasm understanding, and constructs a multi-view representation through the tensor product of such three types of cues. We first ask the LLM to extract linguistic, contextual, and emotional cues respectively via a simple prompt. Taking linguistic cue extraction as an example:
Instruction: Please extract the linguistic cues from the input sentence for sarcasm detection, such as keywords, rhetorical devices, punctuation and language style. Input: [$\mathcal{X}$]
We take the outputs of the LLM’s final hidden layer as the embeddings of the linguistic, contextual and emotional cues, and apply a tensor fusion mechanism to fuse the cues as additional inputs to the sarcasm detection prompt. Inspired by the success of tensor fusion network (TFN) for multi-modal sentiment analysis Zadeh et al. (2017), we apply token-wise tensor fusion to aggregate the cues. In particular, the embeddings are projected on a low-dimensional space, i.e., $\vec{Lin}=\left(e_{1}^{l},e_{2}^{l},...,e_{L}^{l}\right)^{T}$, $\vec{Con}=\left(e_{1}^{c},e_{2}^{c},...,e_{L}^{c}\right)^{T}$, $\vec{Emo}=\left(e_{1}^{e},e_{2}^{e},...,e_{L}^{e}\right)^{T}$. Suppose the LLM has a hidden dimensionality of $d$, fully-connected layers $f_{lin},f_{con},f_{emo}$ are constructed to map the embeddings to dimensionality of $\{d_{l},d_{c},d_{e}\}$, respectively for linguistic, contextual and emotional cues. Then, a tensor product is computed to combine the cues into a high-dimensional representation $\mathcal{Z}=\left(e_{1},e_{2},...,e_{L}\right)^{T}$, where
@@
Consider the information provided in the current cue above. Classify whether the input text is sarcastic or not. If you think the Input text is sarcastic, answer: yes. If you think the Input text is not sarcastic, answer: no. Input: [$\mathcal{X}$]
@@
ToC facilitates deep interactions among these cues, providing a powerful and flexible framework for processing complex linguistic phenomena (see Algorithm 4 in App. A). Notably, as ToC manipulates cues on the vector level via neural structures, it requires access to the LLM structure and calls for supervised training on a collection of labeled samples. During training, the weights of the LLM are frozen, and the linear weights in $f_{lin},f_{con},f_{emo}$ are updated as an adaptation of LLM to the task context.
@@ 4.1 Experiment Setups
MUStARD is compiled from popular TV shows including Friends, The Golden Girls, The Big Bang Theory, etc. It consists of 690 samples total of 3,000 utterances. Each sample is a conversation consisting of several utterances. In this work, we only use the textual information. The statistics for each dataset are shown in Table 2.
Table 2: Dataset statistics.
Dataset Avg. Length #Train #Dev #Test
IAC-V1 68 1,595 80 320
IAC-V2 43 5,216 262 1,042
SemEval 2018 14 3,634 200 784
MUStARD 14 552 - 138
@@
  • PLMs.  (1) RoBERTa Liu et al. (2019), (2) BNS-Net Zhou et al. (2023), (3) DC-Net Liu et al. (2022), (4) QUIET Liu et al. (2023a) and (5) SarcPrompt Liu et al. (2023b) are five SOTA PLMs based approaches for sarcasm detection via pre-trained language modeling and refined representations.
  • Prompt tuning.  (6) IO, (7) CoT Wei et al. (2022) and (8) ToT Yao et al. (2024) are four SOTA prompting approaches by leveraging advanced prompt approaches to enhance LLM’s performance.
  • LLMs.  (9) GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ and (10) LLAMA 3-8B-Instruct [^3] https://llama.meta.com/llama3/ are the strongest general LLMs.
Implementation. We have implemented the prompting methods for GPT-4o and LLaMA3-8B-Instruct, and reported the performance of PLMs in their original papers. The GPT-4o methods are implemented with the official openAI Python API library [^4] https://github.com/openai/openai-python, while the LLaMA methods are implemented based on the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers. All prompting strategies are implemented for GPT-4o and LLaMA3-8B-Instruct except for ToC, which can solely be deployed on open-sourced LLMs. Following previous works in this field, LangChain [^6] https://github.com/langchain-ai/langchain is employed for the implementation of ToT and GoC. For the training of ToC, cross-entropy loss between the output logit and the true label token is computed to update the weights of the fully-connected layers.
Table 3: Performance on four datasets. For LLMs, all strategies but ToC are based on a zero-shot setting.
@@
GoC (Ours) 85.38 68.08 64.97 61.30 74.03 74.02 70.69 69.91 68.33
BoC (Ours) 68.75 67.36 71.35 69.39 62.12 61.85 69.42 68.45 66.79
LLaMA 3-8B-Instruct IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99 44.90
@@
GoC (Ours) 57.10 54.96 42.20 41.61 57.33 57.24 52.77 52.67 51.62
BoC (Ours) 62.50 59.28 62.57 58.11 59.82 58.40 59.71 56.70 58.12
ToC (Ours) 70.31 70.29 79.08 79.07 77.93 76.86 73.33 72.85 74.77
@@ 4.2 Main Results
We report both Accuracy and Macro-F1 results for SarcasmCue and baselines in a zero-shot setting in Table 3, except for ToC which requires supervised training for context adaption.
LLMs do not possess a unique advantage on sarcasm detection. Since sarcasm indicates the manifestation of sentiments and intentions opposite to the literal meaning of the texts, it usually violates logical reasoning pipelines that LLMs are known to excel at Wei et al. (2022). This is empirically validated in the experiment where LLMs are observed to have consistently lower performance over PLMs in terms of average F1 scores across the four datasets. This highlights the need to investigate prompting strategies for adapting LLMs for sarcasm detection, towards which this work has made the first attempt and achieved preliminary success.
Human sarcasm detection does not necessarily follow a step-by-step reasoning process. The comparison between sequential (CoT, CoC, GoC, ToT) and non-sequential (BoC, ToC) prompting strategies fails to provide clear empirical evidences on whether sarcasm detection follows a step-by-step reasoning process. Nevertheless, the results on LLaMA3-8B-Instruct are more indicative to GPT-4o, since the latter has a strong capacity on its own (IO) and does not significantly benefit from any prompting strategies on its top. On LLaMA3-8B-Instruct where in-context learning is necessary for sarcasm detection due to its poor IO performance, non-sequential approaches can apparently offer more benefits over sequential ones, with a remarkable margin consistently present on all four datasets. This seems to support our hypothesize that sarcasm has a non-sequential nature.
@@
SarcasmCue successfully adapts LLMs to sarcasm detection. The proposed prompting strategies in the SarcasmCue framework achieve an overall superior performance to the baseline prompting methods and bring about accuracy increase over the original LLMs in a zero-shot setting. In particular, by explicitly designing the reasoning steps for sarcasm detection, CoC beats CoT by a tremendous margin on GPT-4o, whilst performing in par with CoT on LLaMA3-8B-Instruct, an interesting result that further suggests the non-squential nature of sarcasm detection. By pre-defining the set of cues on 3 main aspects, GoC and BoC manage to guide LLMs to reason along the correct paths, leading to more accuracy judgment of sarcasm than the freestyle thinking in ToT. The proposed trainable neural architecture in ToC achieves an effective tensor fusion of multi-aspect cues for sarcasm detection, pushing the capacity to a comparable level to PLMs without tuning the LLM parameters.
5 Acknowledgments
This work is supported by National Science Foundation of China under grant No. 62006212, Fellowship from the China Postdoctoral Science Foundation (2023M733907), Natural Science Foundation of Hunan Province of China (242300421412).
6 Conclusion

### v2

@@ 3.3 Graph of Cues
The linear structure of CoC restricts it to a single path of reasoning. To fill this gap, we introduce graph of cues, a graph based paradigm that allows LLMs to flexibly choose and weigh multiple cues, unconstrained by the need for unique predecessor nodes (Fig. 2 (b)). GoC frames the problem of sarcasm detection as a search over a graph, and is formulated as a tuple $\left(\mathcal{M},\mathcal{G},\mathcal{E}\right)$, where $\mathcal{M}$ is the cue maker used to define what are the common cues, $\mathcal{G}$ is a graph of “sarcasm detection process”, $\mathcal{E}$ is cue evaluator used to determine which cues to keep selecting.
@@
1. Cue maker. Human sarcasm judgment often relies on the combination and analysis of one or more cues to achieve an accurate understanding. Such cues can be broadly categorized into three types: linguistic cues, contextual cues and emotional cues. Linguistic cues refer to the linguistic features inherent in the text, including keywords, rhetorical devices, punctuation and language style. Contextual cues refer to the environment and background of the text, including topic, cultural background, common knowledge. Emotional cues denote the emotions implied in the text, including emotional words, special symbols (such as emojis) and emotional contrasts. Hence, GoC can obtain 4+3+3=10 cues.
@@
2. Graph construction. In $\mathcal{G}=\left(V,E\right)$, 10 cues are regarded as vertices, constituting the vertex set $V$, the supplement relations across cues are regarded as edges. Given the cue $c_{k}$, the cue evaluator $\mathcal{E}$ considers cue $c_{j}$ to provide the most complementary information to $c_{k}$, which would combine with $c_{k}$ to facilitate a deep understanding of sarcasm.
@@
3. Cue evaluator. We associate $\mathcal{G}$ with LLM detecting sarcasm process. To advance this process, the cue evaluator $\mathcal{E}$ assesses the current progress by asking the LLM whether the cumulative cues obtained thus far are sufficient to yield an accurate judgment. The search goes to an end if a positive answer is returned; otherwise, the detection process proceeds by instructing the LLM to determine which additional cues to select and in what order. In this work, an LLM will act as the cue evaluator, similar to ToT.
@@
We employ a voting strategy to determine the most valuable cue for selection, by deliberately comparing multiple potential cue candidates in a voting prompt, such as:
@@
Given an input text $\mathcal{X}$, the target is to accurately detect sarcasm. Now, we have collected the keyword information as the first step: {keywords}, judge if this provides over 95% confidence for accurate detection. If so, output the result. Otherwise, from the remaining cues {rhetorical devices, punctuation, …}, vote the most valuable one to improve accuracy and confidence for the next step.
@@
This step can be formulated as $\mathcal{E}\left(\mathcal{L}_{\theta}^{GoC},c_{j+1}\right)\sim Vote\left\{\mathcal{L}_{\theta}^{GoC}\left(c_{j+1}|\mathcal{X},c_{1,2,...,j}\right)\right\}_{c_{j+1}\in\{c_{j+1},...,c_{k}\}}$. Until the final judgment is reached, the most valuable cue are always selected in a greedy fashion. Although GoC enables the exploration of many possible paths across the cue graph, its nature remains grounded in a step-by-step reasoning paradigm (see Algorithm 2 in App. A).
@@ 3.4 Bagging of Cues
We relax the assumption that the cues are interrelated in detecting sarcasm. We introduce bagging of cues, a ensemble learning based paradigm that allows LLMs to independently consider varied combinations of cues without assuming a fixed order or dependency among them (Fig. 2 (c)).
BoC constructs a pool of the pre-defined 10 cues $\mathcal{C}$. From this pool, $\mathcal{T}$ subsets are obtained through $\mathcal{T}$ random samplings, where each subset $\mathcal{S}_{t}$ consists of $q~{}\left(i.e.,1\leq q\leq 10\right)$ cues. BoC thus leverages LLMs to generate $\mathcal{T}$ independent sarcasm predictions $\hat{y}_{t}$ based on the cues of each subset. Finally, such predictions are aggregated using a majority voting mechanism to produce the final result. This approach embraces randomness in cue selection, enhancing the LLM’s ability to explore numerous potential paths. BoC consists of three key steps:
Step 1. Cue subsets construction. A total of $\mathcal{T}$ cue subsets $\mathcal{S}_{t\in[1,2,...,\mathcal{T}]}=\left\{c_{t1},c_{t2},...,c_{tq}\right\}$ are created by randomly sampling without replacement from the complete pool of cues $\mathcal{C}$. Each sampling is independent.
Step 2. LLM prediction. For each subset $\mathcal{S}_{t}$, a LLM $\mathcal{L}_{\theta}^{BoC}$ is used to independently make sarcasm prediction through the comprehensive analysis of the cues in the subset and the input text. This can be conceptually encapsulated as $\hat{y}_{t}\sim\mathcal{L}_{\theta}^{BoC}\left(\hat{y}_{t}|\mathcal{S}_{t},\mathcal{X}\right)$.
Step 3. Prediction aggregation. Such predictions $\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{\mathcal{T}}\}$ are then combined using majority voting to yield the final prediction: $Y$.
BoC does not follow the step-by-step reasoning paradigm for sarcasm detection (see Algorithm 3 in App. A.)
@@ 3.5 Tensor of Cues
CoC and GoC methods mainly handle low-order interactions between cues, while BoC assumes cues are independent. To capture high-order interactions among cues, we introduce tensor of cues, a stereo paradigm that allows LLMs to amalgamate three types of cues ($viz.$ linguistic, contextual and emotional cues) into a high-dimensional representation. (Fig. 2 (d)).
ToC treats each type of cues as an independent, orthogonal view for sarcasm understanding, and constructs a multi-view representation through the tensor product of such three types of cues. We first ask the LLM to extract linguistic, contextual, and emotional cues respectively via a simple prompt. For example:
Extract the linguistic cues from the input sentence for sarcasm detection, such as keywords, rhetorical devices, punctuation and language style.
We take the outputs of the LLM’s final hidden layer as the embeddings of the linguistic, contextual and emotional cues, and apply a tensor fusion mechanism to fuse the cues as additional inputs to the sarcasm detection prompt. Inspired by the success of tensor fusion network (TFN) for multi-modal sentiment analysis Zadeh et al. (2017), we apply token-wise tensor fusion to aggregate the cues. In particular, the embeddings are projected on a low-dimensional space via the fully-connected layers, i.e., $\vec{Lin}=\left(e_{1}^{l},e_{2}^{l},...,e_{L}^{l}\right)^{T}$, $\vec{Con}=\left(e_{1}^{c},e_{2}^{c},...,e_{L}^{c}\right)^{T}$, $\vec{Emo}=\left(e_{1}^{e},e_{2}^{e},...,e_{L}^{e}\right)^{T}$. Then, a tensor product is computed to combine the cues into a high-dimensional representation $\mathcal{Z}=\left(e_{1},e_{2},...,e_{L}\right)^{T}$, where
@@
Consider the information provided in the current cue above. Classify whether the input text is sarcastic or not. If you think the Input text is sarcastic, answer: yes. If you think the Input text is not sarcastic, answer: no.
@@
ToC facilitates deep interactions among these cues (see Algorithm 4 in App. A). Notably, as ToC manipulates cues on the vector level via neural structures, it requires access to the LLM structure and calls for supervised training on a collection of labeled samples. During training, the weights of the LLM are frozen, and the linear weights in $f_{lin},f_{con},f_{emo}$ are updated as an adaptation of LLM to the task context.
@@ 4.1 Experiment Setups
Datasets. Four benchmarking datasets are selected as the experimental beds, $viz.$ IAC-V1 Lukin and Walker (2013), IAC-V2 Oraby et al. (2016), SemEval 2018 Task 3 Van Hee et al. (2018) and MUStARD Castro et al. (2019). The details and statistics for each dataset are shown in Table 1 in App. B.
@@
  • Prompt tuning.  (1) IO, (2) CoT Wei et al. (2022) and (3) ToT Yao et al. (2024) are three SOTA prompting approaches by leveraging advanced prompt approaches to enhance LLM’s performance.
  • LLMs. We involve four general LLMs in the experiment, including (4) GPT-4o, (5) Claude 3.5 Sonnet, (6) Llama 3-8B and (7) Qwen 2-7B Bai et al. (2023). The first two are non-open-source LLMs while the last two are open-source LLMs. All four LLMs are representative of the strongest capabilities of their kinds.
Implementation. We have implemented the prompting methods for GPT-4o, Claude 3.5 Sonnet, Llama 3-8B and Qwen2-7B. The GPT-4o and Claude 3.5 Sonnet methods are implemented with the respective official Python API library: openAI [^3] https://github.com/openai/openai-python and anthropic [^4] https://github.com/anthropics/anthropic-sdk-python, while the LLaMA and Qwen methods are implemented based on the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers. Further details are presented in App. C.
Table 2: Performance on four datasets. For LLMs, all strategies are based on a zero-shot setting. Blue and purple indicate the best and second-best results for each dataset. $\clubsuit$ represents significance improvement over the best baseline via unpaired t-test (p $<$ 0.05).
@@
GoC (Ours) 65.00 62.91 64.97 61.30 74.03 74.02 70.69 69.91 67.04
BoC (Ours) 68.75 67.36 71.35 69.39 62.12 61.85 69.42 68.45 66.76
Claude 3.5 Sonnet IO 66.56 66.54 76.78 76.62 75.13 75.11 74.78 74.78 73.26
CoT 71.25 71.14 74.66 74.10 71.56 71.47 73.62 73.53 72.56
ToT 63.44 62.48 71.88 71.74 68.62 68.61 58.84 54.46 64.32
[…]
GoC (Ours) 70.94 70.93 74.67 74.18 76.91 76.91 70.00 69.85 72.97
BoC (Ours) 66.88 66.40 73.61 72.82 70.28 70.07 72.61 71.93 70.31
Llama 3-8B IO 55.94 46.40 54.70 43.74 49.36 44.46 54.64 44.99 44.90
@@
GoC (Ours) 57.10 54.96 54.22 53.30 57.33 57.24 52.77 52.67 54.54
BoC (Ours) 62.50 59.28 62.57 58.11 65.94 65.50 59.71 56.70 59.90
ToC (Ours) 62.19 61.78 72.95 72.94 68.88 68.21 61.26 58.03 65.24
Qwen 2-7B IO 56.56 49.32 51.82 38.57 45.15 38.83 54.78 46.17 43.22
CoT 54.69 46.53 52.88 40.12 43.24 35.79 54.93 45.81 42.06
ToT 53.44 43.71 50.29 39.62 44.26 38.12 52.90 44.60 41.51
[…]
GoC (Ours) 55.00 47.35 53.45 42.25 45.03 38.17 54.49 47.49 43.82
BoC (Ours) 52.50 43.78 52.40 40.24 49.87 45.63 54.06 46.11 43.94
ToC (Ours) 71.56 71.56 72.33 71.76 68.88 68.77 65.94 61.46 68.39
@@ 4.2 Main Results
We report both Accuracy and Macro-F1 scores for SarcasmCue and baselines in Table 2.
@@
(1) SarcasmCue consistently outperforms SoTA prompting baselines. The proposed prompting strategies in the SarcasmCue framework achieve an overall superior performance compared to the baselines and consistently push the SoTA by 4.2%, 2.0%, 29.7% and 58.2% on F1 scores across four datasets. In particular, by explicitly designing the reasoning steps for sarcasm detection, CoC beats CoT by a tremendous margin on GPT-4o and Claude 3.5 Sonnet, whilst performing in par with CoT on Llama 3-8B and Qwen 2-7B. By pre-defining the set of cues in three main categories, GoC and BoC effectively guide LLMs to reason along correct paths, leading to more accurate judgments of sarcasm compared to the freestyle thinking in ToT. For example, the best proposed method, CoC (74.74), brings a 2.0% improvement over the best baseline method, IO (73.26). ToC achieves an effective tensor fusion of multi-aspect cues for sarcasm detection, significantly outperforming other baselines. For instance, it exhibits a 29.7% improvement over the best baseline method, ToT (50.31).
(2) Sarcasm detection does not necessarily follow a step-by-step reasoning process. The comparison between sequential (CoT, CoC, GoC, ToT) and non-sequential (BoC, ToC) prompting strategies fails to provide clear empirical evidences on whether sarcasm detection follows a step-by-step reasoning process. Nevertheless, the results on Llama 3-8B are more indicative to GPT-4o and Claude 3.5 Sonnet, since the latter models have strong capabilities on their own (IO) and do not significantly benefit from any prompting strategies. For Llama 3-8B and Qwen 2-7B, non-sequential methods, particularly ToC, show superior performance. In Llama 3-8B, ToC achieves an average F1 score of 65.24%, which is 8.9% higher than the best sequential method (GoC at 54.54%). The difference is even more pronounced on Qwen 2-7B. This seems to support our hypothesize that sarcasm has a non-sequential nature.
Table 3: Ablation study of BoC, GoC and ToC. All strategies are run on a zero-shot setting. The best results for each dataset are colored in blue.
LLMs Method IAC-V1 IAC-V2 SemEval MUStARD Avg. of F1
Claude 3.5 w/o Lin 68.41 75.62 77.42 69.66 72.78
w/o Emo 69.65 74.04 78.70 70.57 73.24
w/o Con 70.53 74.91 76.39 70.11 72.99
GoC 70.93 74.18 76.91 69.85 72.97
w/o Lin 45.89 42.49 47.47 65.33 50.30
w/o Emo 58.00 56.99 56.81 68.84 60.16
w/o Con 61.71 63.70 69.53 74.80 67.44
BoC 66.40 72.82 70.07 71.93 70.31
Llama 3-8B w/o Lin 45.79 51.90 56.01 46.84 50.14
w/o Emo 48.60 49.40 52.38 45.12 48.88
w/o Con 52.51 53.69 52.14 48.28 51.66
GoC 54.96 53.30 57.24 52.67 54.54
w/o Lin 52.71 57.51 57.53 53.06 55.20
w/o Emo 57.33 59.40 62.01 53.06 57.95
w/o Con 56.88 60.36 59.04 52.30 57.15
BoC 59.28 58.11 65.50 56.70 59.90
w/o Lin 53.31 67.05 59.20 48.05 56.90
w/o Emo 57.42 67.08 64.01 52.89 60.35
w/o Con 55.26 71.78 63.93 52.48 60.86
ToC 61.78 72.94 68.21 58.03 65.24
4.3 Ablation Study
Table 3 presents the result of ablation study. w/o Lin, w/o Emo, w/o Con refer to the method where linguistic, emotional and contextual cues are ablated, respectively. To avoid proactive extraction of ablated cues by an LLM, we explicitly “prompt away” the cues in the inputs. An example prompt could be “You can only use the emotional cues and contextual cues, and do not use any linguistic information here” for the w/o Lin case.
The experiment results highlight the following conclusions: (a) the removal of any single type of cue leads to a noticeable drop in performance across all datasets, demonstrating the importance of each type of cue in sarcasm detection; (b) linguistic cues appear to have the most significant impact, as removing them leads to a noticeable decrease in performance across most settings; (c) the absence of contextual cues also affects the performance, but to a lesser extent compared to linguistic cues.
4.4 Zero-shot v/s Few-shot Prompting
Since the above experiments are mainly based on a zero-shot setting, we are curious of whether the conclusions also apply in a few-shot scenario. Therefore, we perform few-shot experiments to evaluate whether the proposed SarcasmCue framework can perform better when a limited number of contextual examples are available. We plot the main results in Fig. 3, we randomly sample $k=\left\{0,1,5,10\right\}$ examples from the training set. Please refer to Table 2, App. D for the full result.
Figure 3: The average Macro-F1 across K-shots for the GPT-4o and Claude 3.5 Sonnet models.
As shown in the plot, the number of demonstrations has a significant impact on the results. For example, CoC appears sensitive to the initial introduction of demonstration examples with a slight descent in performance when only 1 example is provided. However, as the number of shots increases to 5 and 10, the performance progressively improves. This trend underscores the effectiveness of CoC in adapting and refining its approach with more examples. In contrast, BoC demonstrates a consistent improvement in performance as the number of shots increases.
Overall, these results demonstrate the robustness and adaptability of the SarcasmCue framework in zero-shot and few-shot scenarios. The framework can effectively utilize limited contextual examples to further improve sarcasm detection, making it suitable for applications where large annotated datasets are not readily available.
Figure 4: The influence of model scale. The figures in the top and bottom correspond to Qwen and Llama models, respectively.
4.5 Influences of LLM scales
In an attempt to study the influence of different LLM scales, we evaluate the performance of sarcasm detection of Qwen and Llama of varying sizes, see Fig. 4.
The key take-aways are two-fold. First, the efficacy of our prompting methods is amplified with increasing model scale. This aligns closely with the key findings of the CoT method Wei et al. (2022). This occurs because when an LLM is sufficiently large, its capabilities for multi-hop reasoning and understanding language are significantly enhanced. Second, ToC exhibits high sensitivity to model scale, performing significantly better in larger models, making it particularly suitable for larger-scale applications. CoC and GoC demonstrate moderate sensitivity, indicating a balance between performance improvement and scalability. BoC offers robust performance even in smaller models, suggesting its utility in resource-constrained scenarios. Overall, our proposed framework has a high adaptability across various model scales by offering suitable methods. Please see Table 3 and Fig. 1, App. E for the full results.
4.6 Error Analysis
Figure 5: The average error rate of the four prompting methods.
Fig. 5 shows the error rates of failure cases in terms of false negative (FN) and false positive (FP) for all four prompting methods in SarcasmCue. CoC, GoC and BoC exhibit higher false positive rates, indicating an over-detection of sarcasm that could lead to the frequent misclassification of normal statements as sarcastic. In contrast, ToC exhibits the lowest overall error rate and the FP and FN rates are indeed much closer to each other, indicating a balanced performance in detecting both sarcastic and non-sarcastic texts. These insights highlight potential directions for future improvements in sarcasm detection methodologies. The higher false positive rates suggest a need for refining these methods to reduce over-sensitivity and improve discrimination between sarcastic and non-sarcastic texts. The detailed case study is presented in App. F.
4.7 Extension to New Task
To evaluation the generalization capability of SarcasmCue, we apply it to another complex affection understanding task, humor detection. We compare our proposed SarcasmCue (where the backbone is GPT-4o) with two supervised PLMs (MFN Hasan et al. (2021) and SVM+BERT Zhang et al. (2024)) on two benchmarking datasets, CMMA Zhang et al. (2024) and UR-FUNNY-V2 Hasan et al. (2019).
As shown in Table 4, our methods (BoC and CoC) surpass the baseline on CMMA, whilst performing in par to the strongest baselines on the UR-FUNNY-V2 dataset. These results highlight the strong generalizability and versatility of our framework, confirming its potential utility across a wide range of affection understanding tasks.
Table 4: Performance on two humor detection datasets.
Method CMMA UR-FUNNY-V2 Avg. of F1
Acc. Ma-F1 Acc. Ma-F1
MFN - - 64.44 64.12 -
SVM+BERT 55.23 54.08 69.62 69.27 61.68
CoC 78.14 58.60 64.08 60.13 65.24
GoC 79.60 57.42 64.89 61.65 65.89
BoC 75.81 58.58 68.71 66.83 67.48
5 Conclusions